40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 147 of 354  Not logged in ELOG logo
ID Date Author Type Category Subject
  10458   Fri Sep 5 05:32:57 2014 JenneUpdateLSCGreen PDH box boosts

[Rana, Jenne, EricQ]

* Too much gain overall on Yend box, needed attenuator on output to get lock.  Rethought gain allocation.  Resoldered board, installed, Ygreen locks nicely.  Error point and control point spectra, box TF and open loop TF data collected, to be plotted.

* Q replaced the Xend box, with a matching TF.

* Locked both arms individually, Yend has lots of low freq fluctuation, Xend has some.  Can't do out of loop measurement since we're going well beyond the range of the PDH signals (Yarm RIN is between 1/2 and 1.) Plot TRX and TRY spectra with ALS lock vs. IR lock to get an idea of what frequencies we have a problem with.

* Tried comm/diff locking anyway.  Works.  Used cm_up script to get CARM to sqrtInvTrans.  Went to powers of about 0.5 (hard to say really, because of fluctuations), put sine at 611.1 Hz, 200 cts onto ETMs (-1*x, +1*y), looked at TF between ALS diff and AS55Q.  Put that amount into the static power normalization spot for AS55.  In steps of 0.1, reduced ALSdiff input matrix elements and increased AS55->DARM element.  2 (3?) times was able to get to AS55Q for DARM.  Lost lock once unknown reason, while reducing CARM offset.  Lost lock once trying to turn on FM4 LSC boost for DARM.

TRX/TRY spectra:

TRX_TRY_ALSvsIR_4Sept2014.pdf

  10457   Fri Sep 5 04:07:44 2014 ericqUpdateGreen LockingX end uPDH Box Replaced

Just a quick note, plots and data will come tomorrow:

I grabbed an unused uPDH board from the ATF (thanks Zach!), and re-stuffed almost the entire thing to match Jenne's latest schematic for the y end box. I also threw some 22uF caps on the regulators, as Koji did with the previous box, to eliminate some oscillations up in the high 10s of kHz. I replaced the tragedy of a box that I created on Wednesday with this new box. The arm locks pretty stably with the boost on, 30 degrees of phase margin with 10kHz UGF, and locks pretty darn reliably. 

Now we should now have two nicely boosted PDH loops. I'll do a noise/loop breakdown again in the upcoming days. 

  10456   Fri Sep 5 03:36:00 2014 ranaUpdateComputer Scripts / Programsmartian wireless tweak

 I changed the Martian wireless router to use channel 10 instead of something random (as it was). Using the Android app 'Wifi Analyzer' we could see that the usual channels are dominated by FlumeLab and Caltech Beaver.

The range from 9-13 looked clean so we put it up there. Also, the signal strength drops from -45 to -70 dBm as we walk from the BS down to the ends. We need to tweak the router position and orientation to give us another 10 dB so that we can reliably run the laptops at the ends.

  10455   Fri Sep 5 00:56:00 2014 ranaSummaryOptical LeversITM OLs recentered: violations found

I re-centered the ITMX & ITMY Optical lever beams today since they were off. First I aligned the beam into the vacuum so that it went through the center of the on table optics and then tweaked the receiver optics alignment.

There are several bad practices on these which probably makes them drift:

  • plastic bases on some lens mounts
  • some lens mounts are fastened with a single dog instead of two
  • there is no need to use dogs on mounts that have screw holes. Just put the mount so that 2 screws with washers can be used. The placement for these is not so critical.
  • Use less steering mirrors! The ITMY OL path has 5 optics the beam enters the vacuum!!!

According to the datasheets, the laser has a beam diameter of 0.6 mm and a divergence angle of 1.3/2 mrad. So we can just calculate the right lens positions next time and not have to experiment with the whole visible laser lens kit.

For next Wednesday's cleanup, someone should volunteer to make the mounts more stable for the ITMs.

  10454   Thu Sep 4 18:30:13 2014 GabrieleSummaryASCOptimal Gouy phase for POP QPD

 Jenne asked me to simulate the signals on POP QPD when moving different mirrors, as a function of the Gouy phase where the QPD is placed.

I used the opportunity to create a MIST simulation file of the entire 40m interferometer, essentially based on my aLIGO configuration file. I used the recycling cavity lengths obtained from our survey, and other parameters from the wiki page. The configuration file is attached (fortymeters.mist).

Coming back to the main simulation, here is the result, both for the "regular" POP QPD and for a 22MHz demodulated one. The Gouy phase is measured starting from PR2. Cavity mirrors are easily decoupled from PRM in the "regular" QPD. As already demonstrated in a previous simulation, ETMs signals are very small in the 22 MHz QPD. Moreover, it is possible to zero the contribution from ITMs by choosing the right Gouy phase, at the price of a reduction of the PRM signal by a factor of 3-4. Simulation files are attached.

pop_qpd_dc.png

 

Attachment 2: fortymeters.mist
###########################################################################
# Configuration file for full dual recycled 40m interferometer
classname FortyMeters
################################################################ Parameters

# General parameters
const Pin 1             # input power

# Mirror parameters
const T_ITM 0.01384     # ITM transmission [from https://wiki-40m.ligo.caltech.edu/Core_Optics]
... 143 more lines ...
Attachment 3: fortymeters_pop_qpd.mist
###########################################################################
# Configuration file for full dual recycled 40m interferometer
classname FortyMetersPOP_QPD
################################################################ Parameters

# General parameters
const Pin 1             # input power

# Mirror parameters
const T_ITM 0.01384     # ITM transmission [from https://wiki-40m.ligo.caltech.edu/Core_Optics]
... 148 more lines ...
Attachment 4: pop_qpd.m
% compile and create simulation class
clear classes
MIST('fortymeters_pop_qpd.mist');
s = FortyMetersPOP_QPD(4);

% set angular motion of ITMs, ETMs and PRM
s.ETMX.setMotionShape('pitch');
s.ETMY.setMotionShape('pitch');
s.ITMX.setMotionShape('pitch');
s.ITMY.setMotionShape('pitch');
... 47 more lines ...
  10453   Thu Sep 4 18:16:20 2014 ericqUpdateLSCRecycling cavity lengths

Koji correctly points out that I naïvely overlooked various factors. With a similar analysis to the wiki page, I get:

  • Ideal arm length of 37.795 m
  • Ideal PRC length of 6.753 m
  • Ideal SRC length of 5.399 m

This means that:

  • The PRC, measured at 6.759m, is 6mm long. 
  • The SRC, measured at 5.474m, is 7.5 cm long

Next step is to see how this may affect our ability to sense, and thereby control, the SRC when the arms are going. 

MIST simulations and plots are in the attached zip. 

Attachment 1: 2014-09-CavityLengths.zip
  10452   Thu Sep 4 16:45:10 2014 JenneUpdateLSCGreen PDH box boosts

As EricQ mentioned in last night's elog, the modifications were made to the Yend (SN 17) uPDH board.

R31 became 49.9 Ohms, R30 became 45.3kOhm, R24 became 1.02k, R16 became 1k, a new flying resistor is tombstoned up against R24 and connected by purple wire to C6 and it is 20k.  C28 is 183nF and C6 is 100nF.  These numbers were used in Q's simulation last night.

 

 

IMG_1712.JPGIMG_1714.JPG

  10451   Thu Sep 4 10:10:23 2014 KojiUpdateLSCRecycling cavity lengths

Com'on. This is just a 60ppm change of the mod frequency from the nominal. How can it change the recycling cav length by more than a cm?

https://wiki-40m.ligo.caltech.edu/IFO_Modeling/RC_lengths

This describes how the desirable recycling cavity lengths are affected by the phase of the sidebands at non-resonant reflection of the arms.

If we believe these numbers, L_PRC = 6.7538 [m] and L_SRC = 5.39915 [m].

Compare them with the measured numbers

  • Lprc = 6.752 m
  • Lsrc  = 5.474 m

You should definitely run MIST to see what is the optimal length of the RCs, and what is the effect of the given length deviations.

  10450   Thu Sep 4 03:12:55 2014 ericqUpdateLSCGreen PDH box boosts

Jenne made her board modifications, and the measured TF agreed with the design. Alas, the green would not lock to the arm in this state. 

I think that the reason is that the new TF does not have nearly as much low frequency gain as the old one, for a given UGF. Thus, for example, the 1Hz noise due to the pendulum resonance, has 30dB less loop gain suppressing it. 

boostedTF.pdf

 

NEED MORE gain.jpg

 

  10449   Thu Sep 4 01:28:32 2014 ericqUpdateLSCRecycling cavity lengths

 Going off some discussion we had at lunch today, here is my current knowledge of the state of cavity lengths. 

Acknowledging that Koji changed the sideband modulation frequency recently, the ideal cavity lengths are (to the nearest mm):

  • Lprc = c / ( 4 * fmod) = 6.773 m
  • Lsrc = c / ( 5 * fmod) = 5.418 m

We when last hand measured distances, after moving PR2, we found:

  • Lprc  = 6.752 m = 2.1 cm short
  • Lsrc  = 5.474 m = 5.6 cm long. 

However, when I looked at the sideband splitting interferometrically, I found:

  • Lprc = 6.759m = 1.4 cm short

This is only 5mm from the hand measured value, so we can believe that the SRC length is between 5 and 6 cm too long. I'm building a MIST model to try and see what this may entail. 

  10448   Thu Sep 4 00:56:44 2014 JenneUpdateLSCGreen PDH box boosts

Okay, went back to the drawing board with Rana and Koji on PDH box stuff.

Currently (at least for the Yend), in the boost OFF state, we have an overall gain of about 50.  This is crazy big.  Also, the zero in the "transfer function stage" is around 1kHz, however our green cavity pole is (calculated) to be around 20 kHz.  Since these are supposed to cancel but they're not, we have a wide weird flat region in our loop TF.

So.  I calculated the changes to the TF stage that I'll need so that I have an increase of about 20 in DC gain, kept the pole at the same ~20Hz, but moved the zero way out to 18kHz.  I also calculated the changes needed for the integrator stage to make it effective at much higher frequency than it was designed for.  Now the pole is at 75 Hz, and the zero will be at 1.6kHz, and the high frequency gain will stay pretty close to the same with and without the boost.

Planned new TF stage:

TFstage_newDesign_3Sept2014.png

Planned boost stage (with and without boost activated):

BoostNoBoost_newDesign_3Sept2014.png

New boost stage only, so you can see the phase:

BoostOnly_newDesign_3Sept2014.png

The schematic, modified to show my planned changes (which I will put in the DCC after I make the changes):

D0901351-v1_3Sept2014.pdf

  10447   Wed Sep 3 20:38:40 2014 ericqUpdateGreen LockinguPDH Box Checkup

The traces were from the front panel output BNCs, but the VGA preamp exhibited this asymmetric saturation at its output.  

In any case, I tried to replace the Xend box's AD8336 with a new one, and in doing so, did some irreparable damage to the traces on the board  I was not able to get a new AD8336 into the board. There are some ATF ELOGs where Zach found the AD8336 noise to be bad at low frequencies (link), and its form factor is totally unsuitable for any design that may involve hand modification, since it doesn't even have legs, just tiny little pads. I suggest we never use it for anything in the future. 

Instead, I've hacked on a little daughter board with an OP27 as an inverting op-amp with the gain resistor on the front panel as its feedback resistor, which can swing from 0 to x20 gain (the old gain setting was around 15dB=~x6). I've checked out the TF and output noise, and they look ok. The board can output both rails as well. 

I don't really like this as a long term solution, but I didn't want to leave things in a totally broken state when I left for dinner. 

  10446   Wed Sep 3 18:42:43 2014 JenneUpdateLSCGreen PDH box boosts

From EricQ's simulations reported in elog 10390, we want to transition from ALS comm to DC transmission signals around 500 pm.  However, around 100 pm, the DC transmission signals have a sign flip, so we don't want to have the ALS swing that close to the CARM resonance.  So.  We want to be at about 500 pm, and not touch 100 pm.  So, we don't want our peak ALS motion to go beyond ~400 pm.  Which means that we need to have less than about 40 pm in-loop RMS, to avoid hitting 400 pm.  This is an ALS requirement, but since the analog PDH box is what forces the end laser to follow the arm cavity, and thus give us information about the arm length fluctuations, the PDH residual noise is part of our sensor noise for the full ALS.  So, we need to have the PDH in-loop RMS be less than 40 pm, integrated from a few kHz down to at least 30 mHz. Recall that above the ALS UGF (of about 200 Hz), the sensor noise will be suppressed by 1/f, so we should take that into account when we are looking at the PDH error signal, before we calculate the RMS motion.

Q also measured the in-loop error signal with the current Yend PDH box in elog 10430, and it looks like most of the RMS is coming from a few hundred Hz.  I designed a hack to the PDH board boost that has a zero at about 2kHz, and a gain of 30 at DC, so that we will win by squishing all that RMS.  Also, it shouldn't be too aggressive, so we should be able to leave it on all the time, and still acquire lock of the green laser to the arm, without having to do triggering.

The board schematic is at DCC D1400294.  The boost is also called the "integrator stage", although it will no longer be a simple integrator.

EDIT, JCD:  This cartoon is not correct for the non-boosted state, doesn't include effect of R16.

BoostCartoon.pdf

  10445   Wed Sep 3 14:07:18 2014 ranaConfigurationGeneralnetgpibdata is working again now

Quote:

I gave the IPs to the bridges. According lines of /etc/hosts in linux1 were updated.

192.168.113.230 WET54G1
192.168.113.231 WET54G2

 I was going through some old Koji elogs to check them for correctness (as I do weekly). I noticed that back in Dec 2013, he made the above illegal modification of IP numbers. 192.168.113.230 was actually the IP for farfalla. Maybe that's why they were conflicting and farfalla not working and Q observing/imagining wireless GPIB dropouts?

I used the Wiki instructions to update the 2 bind9 files with a new number for farfalla (192.168.113.212) which was previously the number for the long dead op240m. Farfalla is restarted and sort of working. 

  10444   Wed Sep 3 04:17:21 2014 JenneUpdateLSCY green ALS (not PDH) needs investigation

Q put the X PDH box back, so that I could try locking, and remember which end is up after a week away.

I am unable to hold ALS comm/diff for any length of time. Only once today did I hold it through the FM3 boost turn-on.  So, I looked at the individual arms.

Xarm, even though it's the one that Q is seeing this saturation problem with, seems fine. 

Yarm however is having trouble holding lock for more than a few minutes at a time.  The green beam stays locked to the arm for ~infinity, so I'm not so worried about the PDH box right now.  If I look at the error and control points of the ALS digital servo, the Yarm is much more noisy above about 20 Hz.  Something that I might think of for this kind of mismatch at higher frequencies is poorly matched whitening / dewhitening, or none at all for the Yarm, however this doesn't look like that to me.  Based on the shape of the spectra, I don't think that we're running into ADC noise. For this plot, both arms are individually locked with ALS feeding back to the ETM, gain magnitude of 15 (Xarm gets a minus sign because of our temperature / beatnote moving direction convention), FMs 1,2,3,5,6 on.  Something that seems critical for getting the Yarm to have the FM3 boost without losing lock is having the SLOW temperature servos on for a little while so that the PZT output (as monitored on the temp servo screen) for the end lasers fluctuate around zero. Right now, both beatnotes are at about 62MHz, with an amplitude of about -31dBm.

Yarm_noisy_above_20Hz.pdf

I still need to do a somewhat more thorough investigation of what might be causing the Yarm locklosses.  Is the length-to-angle decoupling worse for ETMY than for ETMX?  Am I moving the arm length so far that the PZT can't follow within its actuation limits?  Does the Yend PDH box have a similar saturation to the Xend box, but somehow (a) worse, and (b) not as obvious so we didn't suspect it before? 

I need to put this plot into calibrated units, and also include the low frequency monitor that we have of the PDH error point (all of which are _DQ channels).

Things to do:

* Figure out Xend PDH box saturation issue.  Is Yend seeing same saturation in the variable gain amplifier?  We have 3 spares of these chips in the Plateau Tournant Bleu, if we need them. 

* Check Yarm ALS stability.  (NB:  The arms have been individually locked for the last 15 min or so while I've been writing, so maybe letting the slow servo settle is the key, and this is not something that needs work).

* Get CARM on DC Trans, DARM on AS55Q (after arm powers of about 1).  Can we see good REFL DC dip?  Should we try using just the transmission PD signal as the error signal for the CM board, if we aren't close enough to resonance to use REFL DC?

  10443   Wed Sep 3 00:17:22 2014 ranaUpdateGreen LockinguPDH Box Checkup

What monitor point is being plotted here? Or is it a scope probe output?

If this saturation is in the uPDH-X but not in the uPDH-Y, then just replace the VGA chip. Because these things have fixed attenuation inside, they often can't go the rails even when the chip is new.

In any case, we need to make a fix to get this box on the air in a fixed state before tomorrow evening.

Quote:

I had noticed in the past, that the digital control signal monitor for the X end would saturate well before the ADC should saturate (C1:ALS-X_SLOW_SERVO_IN1, which is from the "output mon" BNC on the box). It turns out that there is some odd saturation happening inside the box itself.

In this scope trace, the servo input is being driven with a 0.02Vpp, 0.1Hz sine wave, gain knob at 1.0. This is bad. 

TEK00006.PNG

Evan and I poked around the board, and discover that for some reason currently unknown to us, the variable gain amplifier (AD8336) can't reach its negative rail, despite the +-12V arriving safely at its power supply pins. 

I also realized that the LF356 in the integrator stage in this box had been replaced with a LT1792 by Kiwamu in ELOG 4373. I've updated my schematic, and will upload both boxes' schematics to the DCC page Jenne created for them. (D1400293 and D1400294)

 

  10442   Tue Sep 2 22:54:27 2014 KojiSummaryLSCphase tracker UGF

FYI and FMI

Phase tracker UGF is  Q_AMP * G * 2 PI / 360 where Q_AMP is the amplitude of the Q_ERR output and G is the gain of the phase tracker.

For example: Q_AMP = 270, G = 4000\ => UGF = 1.9kHz

  10441   Tue Sep 2 20:02:06 2014 ericqUpdateGreen LockinguPDH Box Checkup

I narrowed down the saturation point in the X green PDH box to the preamp inside the AD8336, but there is still no clear answer as to why it's happening. 

As per Jenne's request, I put the X end PDH box back for tonight's work. It locks, but we have an artificially low actuation range. With SR785, I confirmed a PDH UGF around 5k. Higher than that, and I couldn't reliably measure the UGF due to SR560 saturations. The analyzer is not currently in the loop. 

Both arms lock to green, but I haven't looked at beatnotes today. 

  10440   Tue Sep 2 16:22:24 2014 JenneUpdateGeneralGame plan: 2 Sept

Slightly updated Game Plan.  Mostly, Q is continuing to check out the Xend PDH box saturation, and I am thinking on what our requirements are for ALS, and thus for the green PDH boxes.

40mToDoList.pdf

  10439   Fri Aug 29 23:53:35 2014 ericqUpdateGreen LockinguPDH Box Checkup

I had pulled out both X and Y servo boxes for inspection, put the Y box back, soldered in a missing op amp power capacitor on the X end box, and had not yet put back the X end box yet because of the saturation issue I was looking into. Otherwise nothing was changed at the ends; I didn't open the tables at all, or touch laser/SHG settings, just unplugged the servo boxes. 

  10438   Fri Aug 29 17:28:07 2014 ranaUpdateGreen LockinguPDH Box Checkup

 

 I've been having trouble locking the X - green for the past few hours. Has there been some configuration change down there that anyone knows about?

I'm thinking that perhaps I need to replace the SHG crystal or perhaps remove the PZT alignment mirrors perhaps. Another possibility is that the NPRO down there is going bad. I'll start swapping the Y-end NPRO for the X-end one and see if that makes things better.

  10437   Thu Aug 28 17:34:20 2014 ericqUpdateGreen LockinguPDH Box Checkup

I had noticed in the past, that the digital control signal monitor for the X end would saturate well before the ADC should saturate (C1:ALS-X_SLOW_SERVO_IN1, which is from the "output mon" BNC on the box). It turns out that there is some odd saturation happening inside the box itself.

In this scope trace, the servo input is being driven with a 0.02Vpp, 0.1Hz sine wave, gain knob at 1.0. This is bad. 

TEK00006.PNG

Evan and I poked around the board, and discover that for some reason currently unknown to us, the variable gain amplifier (AD8336) can't reach its negative rail, despite the +-12V arriving safely at its power supply pins. 

I also realized that the LF356 in the integrator stage in this box had been replaced with a LT1792 by Kiwamu in ELOG 4373. I've updated my schematic, and will upload both boxes' schematics to the DCC page Jenne created for them. (D1400293 and D1400294)

  10436   Thu Aug 28 11:02:53 2014 SteveUpdateCalibration-RepairSR785 repair

SN 46,795 of 2003 is back.

Attachment 1: 08281401.PDF
08281401.PDF
  10435   Thu Aug 28 08:31:16 2014 SteveUpdateGeneralone good day
Attachment 1: 1goodDay.png
1goodDay.png
Attachment 2: 1gooday.png
1gooday.png
  10434   Thu Aug 28 01:41:03 2014 ericqUpdateLSCPhase Tracker UGF normalization

We want both the X and Y phase trackers to have the same UGF, so that the X and Y ALS signals are subject to the same phase characteristics and can be nicely decoupled into CARM/DARM. 

I've started implementing a simple normalization scheme that Koji suggested, namely, dividing the I output of the phase tracker by a low passed version of the Q output. (Since the I is servoed to zero, the radius of the error signal in the IQ plane is essentially equal to the Q value) I put some simulink logic into the IQLOCK library part that BEAT[XY]_FINE are instances of to switch the normalization on/off, and to protect from divide-by-zeros. I also exposed the switching and FM on the ALS screen.

UGFnorm.png

I then tried using it, to mediocre results. I put a 10mHz LP in the filter module, found a Y-Arm beat, set the phase tracker gain to give me a 2kHz UGF, and then set the gain of the UGH normalization FM to turn the current average Q to unity. 

I then moved the laser temperature around to get different beatnote locations/amplitudes, hoping that the phase tracker UGF would stay the same when the UGH normalization was on.

It did not.

It did, however, correct it in the right direction... more work will be done with this, to try and make it useful. There's also the unfortunate effect that locking/unlocking the green causes erratic phase tracker output, which messes with the input to the normalizing LP filter, so if one were to leave it switched on, wonky stuff would come out. I don't want to go overboard with triggering shenanigans before I even get it working in the first place, though.  

  10433   Wed Aug 27 18:03:47 2014 ericqUpdateGreen LockinguPDH Box Checkup

 Quick post of plots and data; I'll fill in more detail tonight. 

TL;DR: I pulled both green PDH boxes and made LISO models, compared TFs and noise levels. 


Pictures of X and Y boards, respectively

uPDH_X.JPGuPDH_Y.JPG

 


TF comparison to LISO. (Normalized to coincide at 1Hz)

updhTFs.pdf

 


Noise comparison to LISO

updhNoises.pdf

 


To Do:

  • Figure out why TFs were made differently. Check PM response curves of PZTs to see if they are fine, or need tweaking.
  • Make boosts useful. Both are currently integrators with corners under 10Hz, which is already pretty suppressed. 
  • I just noticed that the X board is missing C25, which should be a 1uF cap on the positive power pin of the primary TF stage opamp. This should be inserted. 

All data, EAGLE schematics, LISO source and plots in the attached zip. 

 

Attachment 5: uPDH_Aug27.zip
  10432   Wed Aug 27 09:12:47 2014 KojiUpdateIMCWFS tuneup

I'm sure that the 1~3Hz motion comes from the mirror motion, but not 100% sure what is causing
the broad stochastic noise. If this is the beam jitter, this penetrates to the IFO via the WFS servos.
Is there any way to characterize this noise in order to compare it with the actual (estimated) motion of the mirrors?

  10431   Tue Aug 26 23:46:55 2014 ericqUpdateIMCWFS tuneup

 I decided to see what I could do with the new WFS setup. 

First, I adjusted the WFS digital demod angles. Once I ensured that the static MC alignment and DC alignment onto the WFS was good, I drove MC2 in pitch with the WFS output off. I then did the usual thing of making the Q peak at the excitation frequency go away. Here are the changes:

  • WFS1 Q1: 7 -> -24 (-31)
  • WFS1 Q2: 6.5 -> -9.5 (-16)
  • WFS1 Q3: -6.5 -> -26.5 (-20)
  • WFS1 Q4: 47 -> 30 (-17)
  • WFS2 Q1: -51 -> -39 (-12)
  • WFS2 Q2: -39 -> -21 (-18)
  • WFS2 Q3: -32 -> -20 (-12)
  • WFS2 Q4: -120 -> -108 (-12)

I then drove each MC mirror in pitch and yaw respectively, and measured the TF from excitation to the WFS signal (dB Magnitude, sign):

 

Mirror DoF WFS1 Pitch WFS1 Yaw WFS2 Pitch WFS2 Yaw
MC1 Pit -67.7,+ -81.9,+ -58.9,- -83.7,+
  Yaw -82.5,- -48.7,- -83.7,+ -112.3,-
MC2 Pit -50.4,- -77.1,- -54.2,- -67.9,+
  Yaw -82.1,- -52.9,+ -59.6,- -44.0,-
MC3 Pit -59.7,- -97.3,+ -62.0,+ -83.9,-
  Yaw -78.0,+ -52.9,+ -67.3,+ -51.4,+

 

I looked through some old ELOG's of Suresh's and used similar logic to scripts/MC/WFS/wfsmatrix2.m to generate a new output matrix. (This involves creating a null sensing vector that is orthogonal to the measured ones, and inverting that matrix) 

Old:

Pitch WFS1 WFS2 MC2T   YAW WFS1 WFS2 MC2T
MC1 -1 0.044 0   MC1 -1 -0.294 0
MC2 0.19 1 1   MC2 -0.26 -0.045 -1
MC3 0.5 -0.681 0   MC3 -.9 1 0

 

New:

 

Pitch WFS1 WFS2 MC2T   YAW WFS1 WFS2 MC2T
MC1 0.835 -1 0   MC1 -1 -0.229 0
MC2 -0.948 -0.433 1   MC2 0.317 -1 -1
MC3 -1 0.865 0   MC3 0.743 0.628 0
 

 

I had to flip a gain or two to keep things stable, then measured the WFS error signal spectra to see if this made anything better. The WFS1 spectra look better, but WFS2 not so much. 

newWFSmatrix.pdf

The loops would need a more thorough investigation, but for now, they're at least a little calmer. The MC is stabler than immediately after the upgrade, but there's still room for improvement. 

 

  10430   Tue Aug 26 23:16:49 2014 ericqConfigurationGreen LockingGain changes on Green Y PDH

 Yesterday I measured the spectra and OLTF of the Y-Arm green PDH, after the LO touch-up and 60Hz hunt from last week. I also went to lower frequencies with the SR785, but forgot to take some of the background spectra down there, so I don't have the full breakdown plots yet. Nevertheless, here is the improvement in the PDH error signal:

pdhComparison.pdf

I also measured the OLTF (SR785 injection at the error signal, Auto level ref 5mV at channel 2, 10mV/s source ramping, 50mV max output)

Ytfs.pdf

As you can see, we have tons of phase margin. Flipping the local boost switch had no visible effect on the OLTF; we should change it to something that puts this surplus of phase to good use, and squash the error signal even more. Putting an integrator at 5kHz should still leave about 45 degrees phase margin at 10k. I've started making a LISO model of the PDH board from the DCC drawing, and then I'll inspect the boards individually to make sure I catch the homegrown modifications. 

Data, and code used to generate the plots is attached. 

Attachment 3: newY.zip
  10429   Mon Aug 25 15:49:44 2014 SteveUpdatePSLHEPA turned on

The PSL HEPA was off. It was turned on and it is running at 30VAC now.

  10428   Mon Aug 25 09:56:21 2014 SteveUpdateIMC IMC WFSs upgrade

 

 The Napa earth quake magnitude 6 did not have any effect on the suspensions.

 The Goy phase upgrade was done nicely. The IOO pointing did not change. Credit owned to Nick and Andres.

  IFO is locked right on.

Attachment 1: eq6Mnapa.png
eq6Mnapa.png
  10427   Fri Aug 22 18:05:02 2014 AndresUpdateIMCUpgrade of the IMC WFSs for the reflection

 Upgrade of IMC Reflection Optical Setup

Nick and I upgrade the IMC. We move both WFSs and placed them facing west. When aligning the beam into the WFS, we make sure that the beam were hitting the center of the mirrors and then we placed the lenses in their corresponding position. We used the beam scanner to measure the waist and the waist in the second WFS was bigger than 1mm, and the second WFS was a little bit below than 1mm. We center the beam in the WFSs and in the PD. We did haven't measure whether we have a good Gouy Phase. Below I attached the picture of how the new setup look like.   

 

Attachment 1: ModeCleanerUpgrade.PNG
ModeCleanerUpgrade.PNG
  10426   Fri Aug 22 18:00:08 2014 jamieOmnistructureCDSubuntu12 awgstream installed

I installed awgstream-2.16.14 in /ligo/apps/ubuntu12.  As with all the ubuntu12 "packages", you need to source the ubuntu12 ligoapps environment script:

controls@pianosa|~ > . /ligo/apps/ubuntu12/ligoapps-user-env.sh
controls@pianosa|~ > which awgstream
/ligo/apps/ubuntu12/awgstream-2.16.14/bin/awgstream
controls@pianosa|~ > 

I tested it on the SRM LSC filter bank.  In one terminal I opened the following camonitor on C1:SUS-SRM_LSC_OUTMON.  In another terminal I ran the following:

controls@pianosa|~ > seq 0 .1 16384  | awgstream C1:SUS-SRM_LSC_EXC 16384 -
Channel = C1:SUS-SRM_LSC_EXC
File    = -
Scale   =          1.000000
Start   = 1092790384.000000
controls@pianosa|~ > 

The camonitor output was:

controls@pianosa|~ > camonitor C1:SUS-SRM_LSC_OUTMON
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:44:50.997418 0  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:49.155525 218.8  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:49.393404 628.4  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:49.629822 935.6  
...
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:58.210810 15066.8  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:58.489501 15476.4  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:58.747095 15886  
C1:SUS-SRM_LSC_OUTMON          2014-08-22 17:52:59.011415 0 

In other words, it seems to work.

  10425   Fri Aug 22 15:58:02 2014 SteveSummaryIOOMC WFS activity

 

 

Attachment 1: GoyphaseSet.png
GoyphaseSet.png
  10424   Fri Aug 22 15:11:55 2014 andres, nicolasSummaryIOOMC WFS activity

1. Before doing anything, we centered the IOO QPDs.
2. With the WFS enabled, we offloaded the control signals onto the bias sliders. Then we saved the slider values. The MC LSC diode had a DC value of ~0.5
3. Turned down power with half wave plate before PMC.  Power injected to vacuum ~ 100mW.
4. We did a beam scan of MC REFL, it looks smaller than what Andres predicted based on the MC eigenmode by 10-20%.
5. We made many changes on the table, pictures to be added by Andres.
6. We didn't have the 80% reflector we wanted to increase the WFS power, so it's still a 98%.
6. Beams were aligned on MC REFL PL, camera, beam dumps, WFSs.
7. Clean up
8. PSL power increased to 1.2W, MC locked right away.
9 We didn't change the IOO WFS output matrix, but we changed some signs and gains to make everything stable. MC autolocker brings it back from cold just fine.
10. All time bombs that we've left will be E.Q.'s to clean up. Sorry.\
11. Yay

  10423   Fri Aug 22 13:51:00 2014 JenneUpdateGeneralUpdated game plan

40mToDoList.pdf

  10422   Fri Aug 22 03:55:45 2014 JenneUpdateLSCXarm PDH fine, Yarm PDH/ALS needs work

[Rana, Jenne, EricQ]

We did several things tonight.  First, a list (so I can remember them all), and then some details.

(1) Jiggled ETMY SUS cables, removed kicks.

(2) Locked X and Y ALS, looked at POX, POY as out of loop sensors.

(3) Measured stuff (?) at the Yend.

(4) Reconnected REFL DC to SR560.

(5) Attempted CARM offset reduction.


Item 1:

When Rana and I started locking this evening, we saw (as Q has been witnessing for a while now) the ETMY kick a lot.  However, it seemed to be kicking even more than usual.  Since Q had been down at the end station recabling things, we wondered if a SUS-related cable got bumped.  Rana went down to the end and pushed all the cables into their receptacles.  One of the last sets that he pushed was the satellite box.  We didn't have walkie-talkie communication, but the DC offset of the ETMY oplevs changed just a minute or two before he returned to the control room.  So, we guess that it was the satellite box cables that were loose.  Unfortunately, there is no clear way to strain relieve them, which is why they can so often be troublesome.  Anyhow, the ETMY hasn't kicked since.

Item 2:

We locked the arms with ALS.  We saw that the POX signal was about 20% of the full pk-pk height of the PDH signal, so it's mostly within the linear range, but not entirely.  It is what it is, however, and we took measurements assuming that it's okay.  I calibrated POX by putting an excitation onto ETMX, and matching the height of the peak in POX and BEATX_FINE_PHASE_OUT_HZ.

Q and Rana had also [remembered / put in / something] a digital readback for the end green PDH error point.  Q went down to the end and gave me a number of 2600 Hz/V for the err mon port of the PDH board, which is what is connected to the ADC.  With that and 20/2^16 V/cts, I had a calibration of 0.8 Hz/ct. 

What we see in this plot is that the green end PDH is not the limiting noise for the POX out of loop measurement of the residual arm motion.  Also, in the multi-color metrology paper, Fig 7 (which is posted in the control room), we see at about a little over 1 Hz a ratio of about 4.5 between the residual motion and the AUX PDH error signal.  In today's plot, I see a ratio of about 20.  I infer from this that the green PDH for the Xarm is fine, and that we may want to re-look at the ALS digital loop, but we should leave the X PDH alone.

Here is the Xarm plot:

Xend_ErrorPointMeasurements.pdf

Q took the data for the Yarm plot, so hopefully he can give it to us in the morning.  What we did notice was that the noise was much worse for the Yarm.  This prompted Item 3, measuring the loop.

Item 3:

Q and Rana went down to the Yend and measured some things.  They came back, and said that they hadn't changed anything in analog while they were down there.  One thing that Q did note was that we have almost 90 degrees of phase margin (since it's a 1/f loop), and about 10 dB of gain margin, above the UGF.  So, we're in good shape for being able to try triggering the boost on the PDH box.  Q will give us more notes on this work, as well as plots, in the morning.

Item 4:

At some point, I remembered that Q and Gabriele had repurposed the SR560 that we had been using for the REFLDC input to the common mode board.  So, Q went and put it back, so that REFL DC goes into the SR560, and so does a DAC channel so that we can remotely set the offset.  The A-B output goes to the REFL11I whitening channel, since real REFL11I goes into the input of the CM board.  I think that today, the SR 560 was left at a gain of 1.

Item 5:

We decided to carry on and try to reduce the CARM offset some.  An annoyance is that the Yarm still has pretty significant low-frequency noise, but the idea is that if we can get over to the sqrtInvTrans signals, it will be fine.

So, we didn't get much farther than we had in the past, but it was nice to get there at all again.  I ran the carm_cm_up script (many times).  One of the times, all I wanted to do was see how much I could reduce the CARM offset.  CARM was on sqrtInvTrans, DARM was on ALS diff, and I was able to get the arm powers up to about 2.5.  I don't know why I lost lock.  The sqrtInv signals should be good until at least arm powers of 20 or so. 

I was able to see the REFL DC dip, but only a teensy tiny bit.  It went down by maybe 1 count.  Q suggested looking at how deep it could get while leaving CARM and DARM both on ALS, and setting both offsets to 0.  We were seeing arm flashes of about 50 counts, and REFL DC went from 0 to -800.  So, I wasn't seeing much of a REFL dip, but it was definitely there when I went to arm powers of 2ish.

We tried looking at different sqrtInv options for DARM, and haven't come to any real conclusion.  In the plot below, we are looking at a swept sine between DARM_IN1 (ALSdiff) and either MC_IN1 0.3*(sqrtInvX - sqrtInvY) or SRCL_IN1 (TRX - TRY / sqrt(TRX + TRY) ):

DARM_ALSdiff_vs_sqrtInv.pdf

 


We have a few things to add to the to-do list:

* Put UGF servos for LSC loops in place.

* Implement UGF "servos" (per Koji's suggested method) for phase trackers.

* Write a lockloss script that is run by the ALS watch scripts - print a PDF of error and control signals for every lockloss, and save it somewhere.

* Fix up Ygreen modematching on the PSL table.  The X green spot is quite similar on the camera to the corresponding PSL green spot.  However the Y green spot is not at all the same as its PSL green spot. 

 

  10421   Thu Aug 21 22:10:52 2014 ranaSummaryComputer Scripts / Programsnetwork movements

 To help development of the data visualization project, we've assigned the .101 and .102 IP to DataVis. This is being used by the iMac in the control room via port 8 of the CDS switch near the Blue Plataeu Tournant.

We tried using one of the free ports, but Jamie realized that we had to use one of the already assigned ones due to some 'Smart' switch management software. So for the moment, please leave the iMac alone so that Bill can use it.

  10420   Thu Aug 21 19:04:52 2014 ericqConfigurationGreen LockingGain changes on Green Y PDH

I found that the barrel of one the BNC to BNC connectors used for getting the output of the PDH servo box to the laser controller was touching the ETMY chamber. When I held it away, all of the 60Hz harmonics disappeared from the mixer output spectrum; this was pretty repeatable. This inspired me to replace the refl PD and PZT signal cables (which were 2 and 3 cables stitched together, respectively) with 20' long BNCs. I also cleaned up a lot of the routing of signal and power cables in the little rack, and moved the big T->DC Block->Attenuator combo off of the panel mount, because I didn't like how it was wiggling. It and the summing pomona box are sitting on top of the PDH box and function generator, instead of hanging freely.

All of the 60Hz harmonics were banished afterwards, and the green locked happily. 

This required me touching the Y end table, to remove the old cable and its cable ties, and putting the new one in. I don't think I did anything immediately apparently bad; the green and IR transmissions both are within nominal ranges. 

I haven't had luck measuring the CLG yet, which I wanted to do to get and set the UGF before measuring the noises. However, here is a scope trace of the in-lock error signal, which compares quite favorably to the trace posted in the previous post; the scope indicates that the signal has 1/3 of the RMS that it did before I replaced the cables. 

TEK00005.PNG

I hope to measure up the current status after I get back from dinner. 

 

  10419   Thu Aug 21 15:07:48 2014 SteveUpdateVACRGA scan at day 197

 

 

Attachment 1: RGAscan@d197.png
RGAscan@d197.png
  10418   Thu Aug 21 02:42:17 2014 rana, ericqConfigurationGreen LockingGain changes on Green Y PDH

[rana, ericq]

We spent time trying to relieve the Yend green PDH of it troubles. 

We realized that the mixer in the PDH setup (mini circuits ZAD-8+), wants 7dBm of LO to properly function. However, we use one function generators output, through a splitter, to give signals to the laser PZT and the mixer LO. 

We don't want 7dBm of power hitting the laser PZT, though. The summing node that adds the servo output to the sideband signal was supposedly designed to do some of this attenuation. Rana measured that 10Vpp out of the function generator resulted in 20mVpp on the fast input to the NPRO, after the summing node. Hence, the 0.09V setting was only resulting in something like 0.2mV hitting the PZT. The PZT has something like 30 rad/V PM response, meaning we only had ~0.006 rad of modulation. 

Now, the function generator is set to 2 Vpp, meaning 4 mVpp hitting the PZT, meaning ~0.12 radians of modulation. The mixer is now getting +7dBm on its LO, and the PDH traces look much cleaner. However, the PDH error signal is now something like 100mVpp, which is much bigger than the PDH board is designed for, so there is now a 10dB attenuator between the reflection PD DC block and the RF input to the mixer. 

Here are screenshots of the Inmon channel (which has a gain of ~20) showing a sweep through some PDH signal, and the error signal while in green lock. Huge 60Hz harmonics are still observed. 

TEK00002.PNGTEK00003.PNG

 


Regarding these 60Hz issues, we need to make sure that we remove all situations where long BNCs are chained together with barrel connectors, or Ts are touching other ones. We also should glue or affix the pomona summing box to the shelf, so that its not just laying on the floor.

The concrete next step is to go fiddle with things, and see if we can get the 60Hz noise to go away, then measure the PDH loop and noises again. Hopefully, this should make the ALS much more reliable. 

  10417   Wed Aug 20 21:09:16 2014 ericqUpdateGreen LockingXarm Green PDH

I remeasured all of the noise spectra again today, making sure the input attenuation was as low as it could safely be. I also got a snap of the y green PDH signal; it's fairly larger than I saw the other day, which is good. I used this to calibrate the error signal voltage spectra. 

scopeSweep.jpg

Here are the noise traces for each arm. During these measurements GTRX was about .6, GTRY about 1.0 The Yarm noise doesn't look so good: the error signal is just barely above the mixer+lowpass output noise, and the RMS is plauged by 60Hz lines. (Is this related to what we see in IR TRY sometimes?)

Xspectra.pdfYspectra.pdf

Here are the arms error signals compared directly:

XYcomp.pdf

  10416   Wed Aug 20 18:05:18 2014 JenneUpdateGreen LockingYarm Green PDH - requirement

 

 I calibrated the control signal from Volts to Hz using the rough PZT calibration of 5MHz/V for the Yend NPRO.  

For the error signal, Q said that the Yarm PDH peak-to-peak height was about a factor of 100 smaller than the Xarm, so I used a calibration of 1.9e7 Hz / V.

Then, from Q's Mist simulation including the high Xarm loss, and the plot that he posted in the control room, the CARM linewidth looks like it is about 2pm.  This is the number that I have included on today's plot.  Note though that yesterday I was using a linewidth of about 30pm, which I got from an Optical simulation about a year ago.  I do not know why these numbers come out an order of magnitude different!      The CARM linewidth is actually about 20 pm.  Both Q and I failed at reading log-x plots yesterday.  I have corrected this, and replotted.

Anyhow, here's the Yarm noise spectra calibrated plot:

YPDH_noise.pdf

I have emailed Kiwamu, but haven't heard back from him yet on what the original design considerations were, if he remembered us ever using a boost, etc.  What this looks like to me is that we need to do some serious work to get the noise down.  Maybe fixing the gain peaking and triggering the boost will get us most of the way there?

  10415   Wed Aug 20 16:10:43 2014 ericqUpdateGreen LockingXarm Green PDH

A MIST simulation tells me that the green pdh horn-to-horn displacement is about 1.2nm, or ~18kHz. I used this, along with the scope trace attached to the previous post, to calibrate the mixer output at 193419 Hz per V. (EDIT: I was a little too hasty here. What I'm really after is the slope of the zero crossing, which turns out to be almost exactly twice my earlier naïve estimate. See later post for correct spectra)

For the control signal, I assumed a flat Innolight PZT PM response of 1MHz/V. ( Under 10kHz, it is indeed flat, and this is the region where the control signal is above the servo output noise in yesterday's measurements)

Here are all of the same spectra from last night, with the above calibrations. 

XspectraCombined.pdf

 

Going off Jenne's earlier plot, it looks like the in-loop error signal RMS is ten times bigger than the CARM linewidth. 

  10414   Wed Aug 20 15:31:27 2014 ericqUpdateCOCArm Loss Investigations Continue

 [ericq, Gabriele] 

Summary: After today's meeting, Gabriele and I looked into the arm loss situation, to see if we should really believe the losses that had been suggested by my previous measurements. We made some observations that we're not sure how to explain, and we're thinking about other ways to try and estimate the losses to corroborate previous findings. 


We first looked to see if the ASS had some effective offset, leaving the alignment not quite right. Once ASS'd, we twiddled each arm cavity mirror in pitch and yaw to see if we could achieve higher transmission. We could not, so this suggested that ASS works properly. 

We then looked at potential offsets in the Xarm loop. We found that an input offset of 25 counts increased the transmission, but only very slightly. With this offset adjusted, we confirmed the qualitative observation that locking/unlocking the xarm causes a much bigger change in ASDC than doing the same with the harm.

However, we noted that the ASDC data (which is the DC value of the AS55 RFPD) was quite noisy, hovering around 50 counts. Looking at the c1lsc model, we found that we were looking at direct ADC counts, so the signal conditioning was not so great. We went to the LSC rack and stole the SR560 that had been hooked up as a REFLDC offsetter, and used it to give ASDC a gain of 100, and a LP at 100Hz, since we only care about DC values. We then undid the gain in the input FM; and this calmed the trace down a fair bit. The effects due to each arm locking/unlocking was still consistent with previous observations. 


At this point, we looked at the arm transmission and ASDC signals simultaneously. Normally, when misaligning a cavity, one would expect the reflected power to rise and the transmission to fall.

However, we saw that when misalignment the Yarm in yaw in either direction, or the Xarm in one direction, both the IR transmission and ASDC would fall. This initially made us think of clipping effects. 

So, we checked out the AS beam situation on the AP table. On a card, the beam looks round as we could tell, and the beam spot on AS55 was nice and small. (We tweaked its steering a little bit in pitch to put it at the center of the "falling-off" points) The reflection and transmission falling effect remained. 


At this point, we're not really sure what could be causing this effect. After the reflected beams recombine at the BS, the output path is common, so it's strange that this odd effect would be the same for both arms. 

Lastly, we discussed other ways that we may be able to see if the Xarm really has ~500ppm loss. Since its transmission is ~1.4%, Gabriele estimated that we may be able to see a ~300Hz difference in the arm cavity pole frequency between the two arms, based on the modification of the cavity finesse due to loss. Since we don't currently have the AOM set up to inject intensity noise, we talked about using frequency noise injection to measure the arm cavity poles, though this would be coupled with the IMC pole, but this could hopefully be accounted for.

  10413   Wed Aug 20 04:09:21 2014 ericqUpdateGreen LockingXarm Green PDH

I've made a whole bunch of measurements on the Xarm green situation.

TL;DRs:

  • GTRX was around 0.55 for all of the measurements tonight. 
  • Based on where I saw gain peaking in the CLG, it looked like UGF was 1-2kHz. I cranked the gain to 10kHz, ~20dB gain peaking followed, making it hard to measure. Currently sitting at 5kHz-ish. 
  • Measured CLG with AG4395A, calibrated for injection point response, inferred OLG. 
  • Took various PSDs, still need to calibrate into physically meaningful units. 

Reasonable amounts of time were spent bending the AG4395 to my will; i.e. figuring out the calibration things Jenne and Rana did, finding the right excitation amplitude and profile that would leave the light steadily locked, and finding the right GPIB incantation for getting spectra in PSD units instead of power units. I'm nearing completion of a newer version of AG4395 scripts that have proper units, and pseudo-log spectra (i.e. logarithmically spaced linear sweeps)


Transfer functions

Here is too many traces on one plot showing parts of the OLTF for the x green PDH. One notable omission is the PD response (note to self:check model and bandwidth). The servo oddly seems to have a notch around 100k. My calibration for the CLG injection may not have been perfect, instead of flattening out at 0dB, I had 2dB residual. I tried to correct for it after the fact, assuming that certain regions were truly flat at 0dB, but I want to revisit it to be thorough. I found some old measurements of the Innolight PZT PM response, which claims to be in rad/V, and have included that on the plot. 

Xbodes.pdf

In the end, the mixer and PZT response make it look like getting over 10kHz bandwidth may be tough. Even finding a good higher modulation frequency to be able to scoot the LP up would leave us with the sharp slope in the PZT phase loss, and could cause bad gain peaking. Maybe it's worth thinking about a faster way of modulating the green light?


Noise Spectra

Tomorrow morning, I'll calibrate all the noise spectra I have into real units. These include:

  • In loop error signal and control signal spectra
  • Mixer output spectrum when PD is dark, and when mixer input is terminated
  • Servo out spectrum when PD is dark, and when servo input is terminated

However, looking at the floors, it occurs to me that I may have left the attenuation on the input too high, in an effort to protect the input the PDH box, which rails all the time when not locked to a 00 mode, sometimes even with the input terminated or open. It's kind of a pain that the agilent makes it really hard to see the data when you're in V/rtHz mode, because I should've caught this while measuring :/

I used a scope to capture a pdh signal happening, which will let me transform the mixer output into cavity motion. The control signal goes to the innolight PZT with a ~1MHz/V factor. Here are the uncalibrated plots, for now. 

scopeSweep.jpg

Xspectra.pdf

 

  10412   Wed Aug 20 02:38:41 2014 JenneUpdateGreen LockingYarm Green PDH - requirement

Quote:

* (JCD) Think about this box's purpose in life.  What kind of gain do we need?  Do we need more / less than we're currently getting? NPRO freq noise is 1/f and is 10kHz/rtHz at 1Hz (this is from a plot of an iLIGO NPRO from Rana's thesis, but it's probably similar). Talk to Kiwamu; the noise budget in the paper seems to indicate that we had some kind of boost on or something.  Also, if we need much more gain than we already have, we'll definitely need a different box, maybe the PDH2 box that they have over in WBridge.

It's not so impressive yet, but here's a plot that shows (a) Rana's guess for laser frequency noise, (b) The inferred in-loop version of that noise, (c) The CARM linewidth FWHM, translated to Hz.

For (b), I take the loop that Rana and I measured last night, and I assumed that it continued on forever as 1/f toward low frequency.  Then I do 1/(1+G) to get the closed loop version of the loop (which is a measurement with an artificial line tacked on the end), and multiply this with the laser freq noise, which is also totally artificial.

For (c), I do df/f = dL/L, with f = c/lambda_green, since the rest of the plot is meant to be in green frequency units.

This is my beginnings of trying to come up with a requirement for our green PDH boxes.  We weren't very clear in the MultiColor paper about the nitty-gritty details (obviously), but then Kiwamu didn't expand on those details in his thesis either.  He talks a lot more about the design considerations for the digital ALS loop, which isn't what I want today.  I will send him an email to see if he had any notes that didn't make it into his thesis.

NoiseConsideration.pdf

  10411   Tue Aug 19 23:11:15 2014 JenneUpdateGreen LockingYarm Green PDH

 

 Here is a plot of last night's data with both the control and the error point on the same plot, in Volts.  Q is still working, so I don't have a calibration number yet to get these to Hz.

Note in the control spectrum that we have very significant 60Hz lines.  

ErrAndCtrlSpectra_VoltsPerRtHz.png

EDIT:  I also added a new branch to the DCC Document Tree, and 2 leafs (one for each end).  Here's the ALS PDH servo branch: E1400350

  10410   Tue Aug 19 21:40:44 2014 AndresUpdateIMCNew Optical Setup for the IMC

IMC Calculation and Setup

I have been working in the calculation for improving the Gouy Phase separation between the WFSs. I tried different possible setup, but the three big constrains in choosing a good optical table setup are to have a Waist size that range from 1mm-2mm, the Gouy Phase  between the WFSs have to be greater than 75 degrees and there has to be a steering mirror before each WFS. I will be showing the best calculation because that calculation complies with Rana request of having both WFSs facing west and having the shortest beam path. I approximate the distances by measuring with a tape the distance where the current optics are located and by looking at the picture that I took I approximated the distance where the lenses will be placed. I'm using a la mode for calculating the gouy phase different. I attached a picture of the current optical table setup that we have. Using a la mode, I found that the current gouy phase that we have is 49.6750 degrees.

Now, for the new setup, a run a la mode and found a Gouy phase of 89.3728 degrees. I have to create a two independent beam path: one for the WFS1 and another one for WFS2. The reason for this is that a la mode place everything in one dimension so and since the WFS1 will have a divergence lens in order to increase the waist size, and since that lens should not be interacting with the waist size in the WFS2. We need two beam path for each WFS.  A la mode give us the following solution:

For the beam path of the WFS1

    label                z (m)           type             parameters        
    -----                  -----              ----             ----------        
    MC1                   0              flat mirror          none:           
    MC3                   0.1753     flat mirror          none:           
    MC2                   13.4587   curved mirror    ROC: 17.8700 (m)     
    Lens1                 29.3705   lens                  focalLength: 1.0201 (m)
    BS2                    29.9475   flat mirror          none:           
    First Mirror         30.0237   flat mirror          none:           
    Lens3                30.2000    lens                  focalLength: -0.100 (m)
    WFS1                30.4809    flat mirror         none: 

For the beam path of the WFS2

    label                   z (m)             type             parameters        
    -----                    -----                 ----             ----------        
    MC1                    0               flat mirror          none:           
    MC3                    0.1753      flat mirror          none:           
    MC2                    13.4587    curved mirror    ROC: 17.8700 (m)     
    Lens1                  29.3705    lens                   focalLength: 1.0201 (m)
    BS2                     29.9475    flat mirror          none:           
    Second Mirror    30.2650     flat mirror          none:           
    Lens2                 30.4809     lens                  focalLength: -0.075 (m)
    Third Mirror        30.5698     flat mirror          none:           
    WFS2                30.6968      flat mirror          none:  

I attached bellow how the new setup should look like in the second picture and also I include and attachment of the a la mode code.

 I used Mist to be able to see the read out that we get in the WFSs that take the Mode Cleaner Reflection and the QPD that take the transmitted from MC2. In the following, plots I'm misaligned the each mirrors: MC1, MC2 and MC3. The misalignment are in Yaw and Pitch. I'm dividing the WFSs reading by the total power reflect power, and I'm dividing the QPD for the MC2 transmission by the total transmitted power. In my Mist model, I have a laser of 1W and my EOM is modulated at 30MHz instead of 29.5MHz and the modulation depth was calculating by measuring the applied voltage using and Spectrum analyzer. I using Kiwamu measurement of modulation depth efficiency vs the applied voltage, https://dcc.ligo.org/DocDB/0010/G1000297/001/G1000297-v1.pdf,  I got a modulation depth of 0.6 mrad. I put this modulation depth and I got the following plots: The fourth and fifth attachment are for the current optical setup that we have. The sixth and seventh attachment is for the new optical setup. The eighth attachment is showing the mode cleaner cavity resonating. The last attachment contains the plots of WFS1 vs WFS2, MC2_QPD vs WFS1, MC2_QPD vs WFS3 for each mirror misaligned. The last two attachment are the MIST code for the calculation.

We have all the lenses that we need. I checked it last Friday and if everything is good we will be ready to do the new upgrade this coming Friday. For increasing the power, I check and we have different BS so we can just switch from the current setup the BS. Can you let me know if this setup look good or if I need to chance the setup? I would really love to do this upgrade before I leave.

 

 

 

 

 

 

Attachment 1: ModeCleanerSetup.PNG
ModeCleanerSetup.PNG
Attachment 2: NewOpticalTableSetupForTheModeCleaner.PNG
NewOpticalTableSetupForTheModeCleaner.PNG
Attachment 3: ReduceWFSPathWorkingOn.m.zip
Attachment 4: MIST_WFSsAndQPDReadingForYaw.png
MIST_WFSsAndQPDReadingForYaw.png
Attachment 5: MIST_WFSsAndQPDReadingForPitch.png
MIST_WFSsAndQPDReadingForPitch.png
Attachment 6: MIST_WFSsAndQPDReadingForYawNewSetup.png
MIST_WFSsAndQPDReadingForYawNewSetup.png
Attachment 7: MIST_WFSsAndQPDReadingForPitchNewSetup.png
MIST_WFSsAndQPDReadingForPitchNewSetup.png
Attachment 8: MISTResonanceCavityReflectionAndTransmissionNewSetup.png
MISTResonanceCavityReflectionAndTransmissionNewSetup.png
Attachment 9: 2Dplots.zip
Attachment 10: ModeCleanerCurrentOpticalTableMIST.zip
Attachment 11: ModeCleanerNewSetupMIST.zip
  10409   Tue Aug 19 18:32:40 2014 ericqUpdateGreen LockingYarm Green PDH

Heading to dinner, going to come back for more green fun, but here's a quick update:

Xarm Peak-to-Peak of the PDH signal in the mixer output is about 70mV when GTRX was about 0.4. The sideband-generating function generator has an output of 2V (forgot to note rms or pp)

Yarm Peak-to-Peak of the PDH signal in the mixer output is about 640uV when GTRX was about 0.71. The sideband-generating function generator has an output of 0.091V (forgot to note rms or pp)

The Yarm signal thus correspondingly has a waaay noisier trace. I would've had scope plots to show here, but the scope freaked out about how large my USB drive capacity was and refused to talk to it >:|

This suggests to me that our modulation depth for the Yarm may be much too small, and may be part of our problems with it. 

ELOG V3.1.3-